OpenStudy (anonymous):

(a) If A is the area of a circle with radius r and the circle expands as time passes, find dA/dt in terms of dr/dt. Below, use "s" for dr/dt. (b) Suppose oil spills from a ruptured tanker and spreads in a circular pattern. If the radius of the oil spill increases at a constant rate of 1m/s, how fast must the area of the spill increasing when the radius is 27m?

OpenStudy (anonymous):

i found the answer for a) 2*pi*r*s but for b i keep getting the wrong answer i keep getting 169.65

OpenStudy (phi):

maybe they want the answer in terms of pi? \[ 54\pi \frac{m^2}{sec} \]

OpenStudy (anonymous):

ok ill try that

OpenStudy (anonymous):

that worked how did you get your answer in pi

OpenStudy (phi):

instead of multiplying out using pi= 3.14159... you leave it pi so 2*pi*27*1 becomes 54*pi or \(54\pi\)

OpenStudy (anonymous):

oh ok that makes sense that thanks for explaining